# Universal Time-Series Representation Learning: A Survey

### v1

@@ 6.1. Evaluation Procedure
Let $\mathcal{D}=\{\mathbf{X}_{i},\mathbf{y}_{i}\}_{i=1}^{N}$ the downstream data. We then compare the encoders $f_{e}$ by using the task-specific evaluation metrics of the downstream task. The evaluation procedures are as follows.
@@ 6.2.2. Classification and Clustering
As the classification and clustering tasks both aim to identify the real category to which a time-series sample belongs, existing studies usually use the same set of benchmark datasets to evaluate the model performance. Curated benchmarks comprising heterogeneous time series from various application domains, such as UCR (Dau et al., 2019) and UEA (Bagnall et al., 2018), are the most widely used because they can provide a comprehensive evaluation regarding the generalization of the model being evaluated. Besides, many researchers also use human activity (e.g., HAR (Anguita et al., 2013)) and health (e.g., PhysioNet Sepsis (Reyna et al., 2020)) related datasets due to their practicality for real-world applications. We recommend referring to Table 3 for the exhaustive list (including audio and video modalities) of datasets for the classification and clustering tasks.
@@ 6.2.3. Regression
Compared to the forecasting and classification tasks, time-series regression, particularly with deep learning, remains relatively underexplored. Only a handful of public benchmark datasets (e.g., heart rate monitoring data (Demirel and Holz, 2023) and air quality (Zhang et al., 2023c)) exist. The TSER archive, introduced by Tan et al. (2021b), is the most comprehensive benchmark for time-series regression. Like forecasting and imputation tasks, the metrics for regression are MSE and MAE. Additional metrics, such as root mean squared error (RMSE) and R-squared ($R^{2}$), are also commonly used.
@@ 6.2.4. Segmentation
Likewise, time-series segmentation with deep learning is also relatively underexplored. There are two standard curated benchmarks: UTSA (Gharghabi et al., 2017) and TSSB (Ermshaus et al., 2023). To assess the segmentation performance, F1 and covering scores are typically used. The F1 score emphasizes the importance of detecting the correct timestamps at which the underlying process changes. In contrast, the covering score focuses on splitting a time series into homogeneous segments and reports a measure for the overlaps of predicted versus labeled segments.
@@ 6.2.5. Anomaly Detection
Anomaly detection is one of the most popular research topics in time series. There are several benchmarks publicly available, as listed in Table 3. However, as argued by recent studies (Wu and Keogh, 2021; Wagner et al., 2023), most existing benchmarks are deeply flawed and cannot provide a meaningful assessment of the anomaly detection models. Therefore, we recommend using newly proposed datasets, such as ASD (Li et al., 2021), TimeSeAD (Wagner et al., 2023), and TSB-UAD (Paparrizos et al., 2022a).
Concerning the evaluation metrics, point-adjust F1 score (Xu et al., 2018) is the most widely used metric for time-series anomaly detection. Nevertheless, this metric is also found to have an overestimation problem, which cannot give a reliable performance evaluation. Accordingly, recent studies (Yang et al., 2023b; Nam et al., 2023; Sylligardos et al., 2023) have started to adopt more robust evaluation metrics, e.g., VUS (Paparrizos et al., 2022b), PA%K (Kim et al., 2022), and eTaPR (Hwang et al., 2022).
@@ 6.2.6. Retrieval
Although we find that a few studies of the reviewed articles use particular datasets for the retrieval task (e.g., EK-100 (Damen et al., 2022), Howto100M (Miech et al., 2019), and MUSIC (Zhao et al., 2018)), the task itself can be evaluated with any benchmark dataset because it is basically based on arbitrary query time series. For time-series retrieval (Chen et al., 2021b), benchmark datasets for classification (e.g., UCR) are commonly used. For the evaluation, the top-$k$ recall rate (higher is better) is the standard metric to examine the overlap percentage of the top-$k$ results and the ground truth. $k$ is usually set to 5, 10, and 20.
@@
Waveform 76,567 2 Medical Data time series (Zhang et al., 2023e)
@@
AD 1,527,552 16 Medical Data time series (Wang et al., 2023b)
PTB 18,711,000 15 Medical Data time series (Wang et al., 2023b)
TDBrain 3,035,136 33 Brain Activity time series (Wang et al., 2023b)
@@
Something-Something 2,650,164 $84\times 84$ Human Activity video (Kim et al., 2023b)
Regression TSER Archive $19\times T$ $V$ Multiple time series (Tan et al., 2021b)
@@
AEPD 19,735 29 Appliances Energy time series (Zhang et al., 2023c)
BMAD 420,768 6 Air Quality time series (Zhang et al., 2023c)
SML2010 4,137 18 Smart home time series (Zhang et al., 2023c)
@@
TSB-UAD $1,980\times T$ 1 Multiple time series (Paparrizos et al., 2022a)
@@
6.3. Additional Metric for Inherent Representation Quality
7. Open Challenges and Future Research Directions
In this section, we discuss open challenges around time-series analysis and outline promising future research directions that have the potential to enhance the existing universal time-series representation learning methods.
7.1. Time-Series Annotation and Active Learning
The complexity and length of time-series data significantly increase the cost of manual labeling. Obtaining annotated time series is challenging due to the domain-specific nature and lack of publicly accessible sources. Even domain experts may struggle to provide consistent annotations because time-series shapes can be perceived differently among annotators. For instance, datasets like Sleep-EDF (Goldberger et al., 2000), containing lengthy EEG recordings, require doctors to identify anomalies with a consensus among multiple experts for each annotation. This level of expertise is also demanded by applications such as anomaly detection in smart factory sensors, making the process time-consuming and labor-intensive.
Efforts within the deep learning community have aimed to address label sparsity, notably through active learning (Shin et al., 2021; Rana and Rawat, 2022). This approach minimizes labeling costs by selecting the most informative unlabeled instances and presenting them to an expert (oracle) for annotation. Active learning involves developing a query strategy that prioritizes which data points to label next based on predefined criteria, e.g., informativeness and diversity. The process iterates by training a model with the initial labels, selecting the most informative or diverse unlabeled data points for labeling, retraining the model, and repeating this cycle until a certain performance threshold (e.g., annotation budget) is reached. Therefore, developing an effective annotation process with active learning for time-series representation learning, enabling a more robust supervision signal, is expected to be a promising research direction.
7.2. Distribution Shifts and Adaptation
Distribution shifts between training and testing data are one of the factors causing the trained model to exhibit low performance during the test phase. When a time series model undergoes continuous testing, especially during the model deployment, various forms of concept drift, such as sudden, gradual, incremental, and recurrent drifts, may occur due to the accumulation of unseen time-series data (Agrahari and Singh, 2022). A domain shift problem may also arise due to differences between the source domain used in the training phase and the target domain used in the test phase.
Given that distribution shifts resulting from concept drift and domain shift are the factors that degrade model performance, previous studies have focused on concept drift adaptation and domain adaptation to address these shifts in specific downstream tasks, such as classification, anomaly detection, and forecasting (Yuan et al., 2022; Ragab et al., 2023; Jin et al., 2022). Addressing distribution shifts in the test phase is also crucial for learning representations for various downstream environments. Therefore, researchers should consider future directions to develop distribution shift adaptations for universal representation learning with discrepancy-based and adversarial methods, for example.
7.3. Reliable Data Augmentation
Determining the appropriate type and degree of data augmentation is essential, as minor changes can significantly alter the original characteristics of time series. Various techniques, including jittering, shifting, scaling, and warping, have been used for time-series representation learning. However, the reliability of these methods has not been fully explored. With the advent of new learning paradigms, such as contrastive learning, the importance of data augmentation increases. Data augmentation serves not only to expand the size of datasets but also to provide diverse class-invariant features. Unfortunately, most current approaches focus more on the evolving role of data augmentation, often overlooking its fundamental task to maintain the integrity of the original data characteristics. Recent studies (Luo et al., 2023; Demirel and Holz, 2023) have started to address the reliability of data augmentation and develop new techniques or adaptive strategies based on more reliable criteria. Despite these advancements, many studies still rely on empirical methods for determining data augmentation. Therefore, a promising direction is devising an evaluation method for time-series data augmentation to estimate the reliability and effectiveness of selecting the potential optimal augmentation strategy.
7.4. Neural Architecture Search
Making efficient deep learning models for universal representation learning is challenging because it laboriously relies on tedious manual trial-and-error to design the network architectures and select corresponding hyperparameters relevant to the three design elements we discussed. The inefficient hand-crafted design will inevitably involve human bias, leading to sub-optimal architectures (Zhang et al., 2023f). Accordingly, recent endeavors employ NAS to discover an optimal architecture by automatically designing neural architectures and network hyperparameters, e.g., the number of layers, network type, embedding dimension, and hidden dimension. These configurations significantly affect the learned representation quality and the downstream task performance. Even though NAS has demonstrated its successes in diverse tasks for computer vision and natural language processing (White et al., 2023), NAS for time series is still underexplored. There are only NAS methods for a specific time-series analysis task, such as forecasting (Shah et al., 2021; Jie and Gao, 2021; Wang et al., 2022; Deng et al., 2022; Lai et al., 2023) and classification (Rakhshani et al., 2020; Xiao et al., 2021; Ren et al., 2022), thus having very limited generalizability. Automatically discovering neural architectures for universal time-series representation learning that perform well across downstream tasks is an important future direction, especially for industry-scale time series having high dimensions and large volumes newly generated every day.
7.5. Large Language and Foundation Models
7.6. Representations of Irregularly-Sampled Time Series
Irregular time series naturally appear in real-world scenarios across various fields, including finance, healthcare, and environmental observation. This irregularity usually stems from sporadic observations and system failures. However, managing irregularly-sampled time series effectively has proven challenging because existing deep learning models are devised initially for regularly-spaced natural language data. Early attempts like GRU-D (Che et al., 2018) modify existing deep models to handle irregular time series. Another line of work uses Neural ODEs (Chen et al., 2018) that learn continuous vector space to represent irregular time series. While the concept of neural differential equations (NDEs) aligns well with irregular time series, these methods still have limitations. First, the training time of NDEs may be instable due to the use of the Runge-Kutta method for integration calculations. Second, the requirement that the hidden state of NDEs should be differentiable limits their ability to represent high-frequency time series. Compared to NDEs, approaches modifying attention (Shukla and Marlin, 2021) or employing interpolation (Shukla and Marlin, 2019) to address irregular time series also show promising results.
In addition to the model-centric approaches, we expect a potential research direction for irregular time series to be a data-centric approach. Unlike regular time-series datasets, irregular time series can contain diverse causes of irregularities often overlooked by previous studies. For instance, medical data may exhibit irregularities from sensor malfunctions, while financial data (e.g., credit card transactions) may contain irregularities due to its sporadic event-based nature. By integrating the understanding of the causes of irregularities into the learning process, we can obtain more precise representations of irregular time series.
7.7. Multi-Modal and Multi-View Representation Learning
Vision-language models such as CLIP (Radford et al., 2021) and ALIGN (Jia et al., 2021) have recently shown remarkable performance in zero-shot learning and fine-tuning for various vision-related downstream tasks by utilizing the semantics of human languages. If time-series data entail semantics understandable to humans, we can make multi-view representations of time series by using human languages as an additional modality to annotate them. As a result, the learned representations from time series-language representation learning will become more expressive and fine-grained in semantics. For example, human activity data (Chen et al., 2021a) or time-series anomalies can be described in human languages based on human movement and the context of abnormal situations. However, unlike video representation learning, which can leverage a vision-language pre-trained model for the annotation, time-series data annotation with human language can be laborious and highly domain-specific. Thus, building a large multi-modal time series-text dataset will be a promising future direction.
@@ 8. Conclusions
This article introduces a universal time-series representation learning research and its importance for downstream time-series analysis. We present a comprehensive and up-to-date literature review of universal representation learning for time series by categorizing the recent advancements from design perspectives. Our main goal is to answer how each fundamental design element—neural architectures, learning objectives, and training data—of state-of-the-art time-series representation learning methods contributes to the improvement of the learned representation quality, resulting in a novel structured taxonomy with fifteen subcategories. Although most state-of-the-art studies consider all design elements in their methods, only one or two elements are newly proposed. Given the current review of the selected studies, we find that decomposition and transformation methods and sample selection techniques in the data-centric approaches are still underexplored. In addition, we provide a practical guideline about standard experimental setups and widely used time-series datasets for particular downstream tasks, together with discussions on various open challenges and future research directions related to time-series representation learning. Ultimately, we expect this survey to be a valuable resource for practitioners and researchers interested in a multi-faceted understanding of the universal representation learning methods for time series.

### v3

@@ 6.1. Evaluation Procedure
Let $\mathcal{D}=\{(\mathbf{X}_{i},\mathbf{y}_{i})\}_{i=1}^{N}$ denote the downstream data. We then compare the encoders $f_{e}$ by using the task-specific evaluation metrics of the downstream task. The evaluation procedures are as follows.
@@ 6.1.2. Fine-Tuning Protocol
End-to-End Protocol. This protocol is a special case for evaluating a representation learning framework, where both $f_{e}$ and $g_{d}$ are trained jointly from scratch for each downstream task without pre-training in (1). Notable examples include TimesNet (Wu et al., 2023), MSD-Mixer (Zhong et al., 2023), and WHEN (Wang et al., 2023c).
@@ 6.2.2. Classification and Clustering
As the classification and clustering tasks both aim to identify the real category to which a time-series sample belongs, existing studies usually use the same set of benchmark datasets to evaluate the model performance. Curated benchmarks comprising heterogeneous time series from various application domains, such as UCR (Dau et al., 2019) and UEA (Bagnall et al., 2018), are the most widely used because they can provide a comprehensive evaluation regarding the generalization of the model being evaluated. Many researchers also use human activity (e.g., HAR (Anguita et al., 2013)) and health (e.g., Sepsis (Reyna et al., 2020)) related datasets due to their practicality for real-world applications.
@@ 6.2.3. Regression
Compared to the forecasting and classification tasks, time-series regression, particularly with DL, remains relatively underexplored. Only a handful of public benchmark datasets (e.g., heart rate monitoring data (Demirel and Holz, 2023) and air quality (Zhang et al., 2023a)) exist. The TSER archive, introduced by Tan et al. (2021), is the most comprehensive benchmark for time-series regression. Like forecasting and imputation tasks, the metrics for regression are MSE and MAE. Additional metrics, such as root mean squared error (RMSE) and R-squared ($R^{2}$), are also commonly used.
@@ 6.2.4. Segmentation
Likewise, time-series segmentation with DL is also relatively underexplored. There are two standard curated benchmarks: UTSA (Gharghabi et al., 2017) and TSSB (Ermshaus et al., 2023). To assess the segmentation performance, F1 and covering scores are typically used. The F1 score emphasizes the importance of detecting the correct timestamps at which the underlying process changes. In contrast, the covering score focuses on splitting a time series into homogeneous segments and reports a measure for the overlaps of predicted versus labeled segments.
@@ 6.2.5. Anomaly Detection
Anomaly detection is one of the most popular research topics in time series. There are several benchmarks publicly available, as listed in Table 3. However, as argued by recent studies (Wu and Keogh, 2021; Wagner et al., 2023), most existing benchmarks are deeply flawed and cannot provide a meaningful assessment of the anomaly detection models. Therefore, we recommend using newly proposed datasets, such as ASD (Li et al., 2021), TimeSeAD (Wagner et al., 2023), and TSB-UAD (Paparrizos et al., 2022b).
Concerning the evaluation metrics, point-adjust F1 score (Xu et al., 2018) is the most widely used metric for time-series anomaly detection. Nevertheless, this metric is also found to have an overestimation problem, which cannot give a reliable performance evaluation. Accordingly, recent studies (Yang et al., 2023b; Nam et al., 2023) have started to adopt more robust evaluation metrics, e.g., VUS (Paparrizos et al., 2022a), PA%K (Kim et al., 2022), and eTaPR (Hwang et al., 2022).
@@ 6.2.6. Retrieval
Although we find that a few studies of the reviewed articles use particular datasets for the retrieval task (e.g., EK-100 (Damen et al., 2022), Howto100M (Miech et al., 2019), and MUSIC (Zhao et al., 2018)), the task itself can be evaluated with any benchmark dataset because it is basically based on arbitrary query time series. For time-series retrieval (Chen et al., 2021a), benchmark datasets for classification (e.g., UCR) are commonly used. For the evaluation, the top-$k$ recall rate (higher is better) is the standard metric to examine the overlap percentage of the top-$k$ results and the ground truth. $k$ is usually set to 5, 10, and 20.
6.3. Additional Metric for Inherent Representation Quality
7. Open Challenges and Future Research Directions
In this section, we discuss open challenges and outline promising future research directions that have the potential to enhance the existing universal time-series representation learning methods.
7.1. Time-Series Active Learning
The complexity and length of time-series data significantly increase the cost of manual labeling. Obtaining annotated time series is challenging due to the domain-specific nature and lack of publicly accessible sources. Even domain experts may struggle to provide consistent annotations because time-series shapes can be perceived differently among annotators. For instance, datasets like Sleep-EDF (PhysioBank, 2000), containing lengthy EEG recordings, require doctors to identify anomalies with a consensus among multiple experts for each annotation. This level of expertise is also demanded by applications such as anomaly detection in smart factory sensors, making the process time-consuming and labor-intensive. Efforts in the DL community have aimed to address label sparsity, notably through active learning (AL) (Shin et al., 2021; Rana and Rawat, 2022). This approach minimizes labeling costs by selecting the most informative unlabeled instances and presenting them to an expert (oracle) for annotation. AL involves developing a query strategy that prioritizes which data points to label next based on predefined criteria, e.g., informativeness and diversity. The process iterates by training a model with the initial labels, selecting the most informative or diverse unlabeled data points for labeling, retraining the model, and repeating this cycle until a certain performance threshold (e.g., annotation budget) is reached. Therefore, developing an effective annotation process with AL for time-series data, enabling a more robust supervision signal, is a promising research direction.
7.2. Distribution Shifts and Adaptation
When a time series model undergoes continuous testing due to the accumulation of time series data over time, various forms of concept drift, such as sudden, gradual, incremental, and recurrent drift, may occur (Agrahari and Singh, 2022). Additionally, a domain shift problem may arise due to differences between the source domain used in the training phase and the target domain used in the test phase. Given that distribution shifts resulting from concept drift and domain shift are factors that degrade model performance, previous studies focus on concept drift adaptation and domain adaptation to address these shifts in specific downstream tasks, such as classification and forecasting (Yuan et al., 2022; Ragab et al., 2023; Jin et al., 2022; Ozyurt et al., 2023). Addressing distribution shifts in the test phase is also crucial for learning representations for various downstream environments. Therefore, researchers should consider future directions to develop distribution shift adaptations for universal representation learning with discrepancy-based and adversarial methods, for example.
7.3. Reliable Data Augmentation
As minor data augmentation can significantly impact time-series properties, determining the appropriate type and degree of augmentation is essential. Various techniques, including jittering, shifting, and warping, have been used for time-series representation learning. However, the reliability of these methods has not been fully explored. Data augmentation becomes more important with new learning paradigms like contrastive learning. It serves not only to expand the size of datasets but also to provide diverse class-invariant features. Unfortunately, current approaches focus more on the evolving role of data augmentation, often overlooking its fundamental task to maintain the integrity of the original data characteristics. Recent studies (Luo et al., 2023; Demirel and Holz, 2023) have focused on improving data augmentation reliability using innovative approaches and adaptive strategies based on more reliable criteria. Many studies still use empirical approaches to determine data augmentation, despite these advances. Therefore, methods for time-series data augmentation that estimates the reliability and efficacy of selecting the optimal augmentation strategy is promising.
7.4. Neural Architecture Search (NAS)
Making efficient deep models for universal representation learning is challenging because it laboriously relies on tedious manual trial-and-error to design the network architectures and select corresponding hyperparameters relevant to the three design elements we discussed. The inefficient hand-crafted design will inevitably involve human bias, leading to sub-optimal architectures (Zhang et al., 2023d). Accordingly, recent endeavors employ NAS to discover an optimal architecture by automatically designing neural architectures and network hyperparameters, e.g., # layers, network type, and embedding dimension. These configurations significantly affect the learned representation quality and the downstream performance. Even though NAS has demonstrated its successes in diverse tasks (White et al., 2023), NAS for time series is still underexplored. There are only NAS methods for a specific time-series analysis task, such as forecasting (Shah et al., 2021; Lai et al., 2023) and classification (Rakhshani et al., 2020; Ren et al., 2022), thus having very limited generalizability. NAS for universal time-series representation learning that perform well across downstream tasks is an important future direction, especially for industry-scale time series having high dimensions and large volumes newly generated every day.
7.5. Large Language and Foundation Models
7.6. Representations of Irregularly-Sampled Time Series (ISTS)
Irregular time series naturally appear in real-world scenarios across various fields, including finance, healthcare, and environmental observation. This irregularity usually stems from sporadic observations and system failures. However, managing ISTS effectively has proven challenging because existing deep learning models are devised initially for regularly-spaced natural language data. Early attempts like GRU-D (Che et al., 2018) modify existing deep models to handle ISTS. Another line of work uses Neural ODEs (Chen et al., 2018) that learn continuous vector space to represent ISTS. While the concept of NDEs aligns well with ISTS, these methods still have limitations. First, the training time of NDEs may be instable due to the use of the Runge-Kutta method for integration calculations. Second, the requirement that the hidden state of NDEs should be differentiable limits their ability to represent high-frequency time series. Compared to NDEs, approaches modifying attention (Shukla and Marlin, 2021) or employing interpolation (Shukla and Marlin, 2019) to address ISTS also show promising results.
Besides the model-centric approaches, we expect a potential research direction for ISTS to be a data-centric approach. Unlike regular time series, ISTS can contain diverse causes of irregularities often overlooked by previous studies. For instance, medical data may exhibit irregularities from sensor malfunctions, while financial data (e.g., credit card transactions) may contain irregularities due to its sporadic event-based nature. By integrating the understanding of the causes of irregularities into the learning process, we can obtain more precise representations of ISTS.
7.7. Multi-Modal and Multi-View Representation Learning
Vision-language models such as CLIP (Radford et al., 2021) and ALIGN (Jia et al., 2021) have recently shown remarkable performance in zero-shot learning and fine-tuning for various vision-related downstream tasks by utilizing the semantics of human languages. If time-series data entail semantics understandable to humans, we can make multi-view representations of time series by using human languages as an additional modality to annotate them. As a result, the learned representations from time series-language representation learning will become more expressive and fine-grained in semantics. For example, human activity data (Chen et al., 2021c) or time-series anomalies can be described in human languages based on human movement and the context of abnormal situations. However, unlike video representation learning, which can leverage a vision-language pre-trained model for the annotation, time-series data annotation with human language can be laborious and highly domain-specific. Thus, building a large multi-modal time series-text dataset will be a promising future direction.
@@ 7.7. Multi-Modal and Multi-View Representation Learning
Waveform 76,567 2 Medical Data time series (Zhang et al., 2023f)
@@
AD 1,527,552 16 Medical Data time series (Wang et al., 2023a)
PTB 18,711,000 15 Medical Data time series (Wang et al., 2023a)
TDBrain 3,035,136 33 Brain Activity time series (Wang et al., 2023a)
@@
Something-Something 2,650,164 $84\times 84$ Human Activity video (Kim et al., 2023c)
Regression TSER Archive $19\times T$ $V$ Multiple time series (Tan et al., 2021)
@@
AEPD 19,735 29 Appliances Energy time series (Zhang et al., 2023a)
BMAD 420,768 6 Air Quality time series (Zhang et al., 2023a)
SML2010 4,137 18 Smart home time series (Zhang et al., 2023a)
@@
TSB-UAD $1,980\times T$ 1 Multiple time series (Paparrizos et al., 2022b)
@@ 8. Conclusions
This article introduces a universal time-series representation learning research and its importance for downstream time-series analysis. We present a comprehensive and up-to-date literature review of universal representation learning for time series by categorizing the recent advancements from design perspectives. Our main goal is to answer how each fundamental design element—training data, neural architectures, and learning objectives—of state-of-the-art time-series representation learning methods contributes to the improvement of the learned representation quality, resulting in a novel structured taxonomy with 26 subcategories. Although most state-of-the-art studies consider all design elements in their methods, only one or two elements are newly proposed. Given the current review of the selected studies, we find that decomposition, transformation, and sample selection methods in the data-centric approaches are still underexplored. In addition, we provide a practical guideline about standard experimental setups and widely used time-series datasets for particular downstream tasks, together with discussions on various open challenges and future research directions related to time-series representation learning. Ultimately, we expect this survey to be a valuable resource for practitioners and researchers interested in a multi-faceted understanding of the universal representation learning methods for time series.